OpenStudy (anonymous):

A car traveling at 60 km/h will skid 30 m when its brakes are locked. If the same car is traveling at 180 km/h, what will be its skidding distance? Show your work

OpenStudy (abhisar):

\(\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\\\color{white}{.}\\\Huge\color{blue}{\mathfrak{~~~~Welcome~to~OpenStudy!~\ddot\smile}}\\\color{white}{.}\\\\\Huge{\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}\color{orange}{\bigstar}\color{red}{\bigstar}\color{blue}{\bigstar}\color{green}{\bigstar}\color{yellow}{\bigstar}}\)

OpenStudy (anonymous):

You sit at the outer rim of a Ferris wheel that rotates at 2 revolutions per minute (RPM). What would your rotational speed be if you were instead clinging to a position halfway from the center to the outer rim?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

At the outer edge of a rotating space habitat, 130 m from its center, the rotational acceleration is g. What is the rotational acceleration at a distance of 65 m from the center of the habitat?

OpenStudy (abhisar):

It would be easier to tutor you if you explained what about the question is giving you trouble.

OpenStudy (abhisar):

Let's begin with ur first question

OpenStudy (anonymous):

okay the problem is I just don't know how to do it. Id appreciate if you explained

OpenStudy (abhisar):

Do you know the equation of motions ?

OpenStudy (abhisar):

V\(^2\) = U\(^2\) + 2aS ?

OpenStudy (anonymous):

no I dont

OpenStudy (abhisar):

ohkay, in that formula V=final speed, U=initial speed, a= accelaration and S=distance covered

OpenStudy (abhisar):

In first case, initial speed = 60km/hr Final speed = 0km/hr Distance covered = 0.03km

OpenStudy (abhisar):

got it ?

OpenStudy (anonymous):

yes

OpenStudy (abhisar):

Now substitute the values and find the missing value of accelaration

OpenStudy (anonymous):

Okay, (0)^2 = (60)^2 + 2a(0.03)

OpenStudy (abhisar):

yes so a = ?

OpenStudy (anonymous):

180?

OpenStudy (abhisar):

a= (-60*60)/0.06

OpenStudy (abhisar):

i.e 60000

OpenStudy (abhisar):

Now for the next case simply plug in the given values U=180km/hr V=0 a=60000 and then find S. This will be the answer

OpenStudy (anonymous):

0 = 180 + 2(60000)(0.03)s right?

OpenStudy (abhisar):

yes, find out s

OpenStudy (anonymous):

0=180+2(1800)s 0=180+3600s 0=3780s s=0?

OpenStudy (abhisar):

a=-60000 u have to put u\(^2\) and not just u, so put (180)\(^2\)

OpenStudy (abhisar):

and why r u putting 0.03 in this case ?

OpenStudy (abhisar):

That was the distance travelled in first case

OpenStudy (anonymous):

0=32400+ 2(-60000)s? I thought I had to put that I got confused

OpenStudy (abhisar):

Here, 0\(^2\) = 180\(^2\) + 2*60000*S

OpenStudy (abhisar):

*-60000

OpenStudy (abhisar):

So, 120000*S = 180 * 180

OpenStudy (abhisar):

S= 32400/120000

OpenStudy (abhisar):

and this will give S= 270 m

OpenStudy (anonymous):

oh okay I get it now I was just confused of why the 0.03 was there.

OpenStudy (abhisar):

30m = 0.03 Km

OpenStudy (abhisar):

It is the distance travelled in first case. got it ?

OpenStudy (anonymous):

yes I got it thanks :)

OpenStudy (anonymous):

could you also help with the other two questions?

OpenStudy (abhisar):

Close this one first and post them as new one !

OpenStudy (abhisar):

Multiple questions in one thread is not allowed

OpenStudy (abhisar):

Did you know you can give medals to people who help you with questions? Just click the \(\color{blue}{\text{blue "best response" button}}\) to show appreciation for the people who help you.

OpenStudy (anonymous):

oh okay ill do that and oh okay ya ill give you a medal.

OpenStudy (abhisar):

No i'm not asking to give me a medal ! It's ur choice. I am only making u aware of this since u look new to the site

OpenStudy (abhisar):

U should not give it to me if u think u r still unsatisfied

OpenStudy (anonymous):

Am satisfied :)